Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 12:59 Sunday 21 February 2021 (16074) Print this report
Comment to Picomotor of PR3_BF_T(Motor2) got stuck (15976)

This is a report about work carried out on Friday 19th of February.

I moved the BF to a position similar to the one in which the BF was before the BF picomotor was moved originally. This position is measured with respect to the BF Damper, which is the BF recoil mass holding LVDTs and coil actuators. The idea is to check whether the alignment of the IM is recovered.

Result: the alignment of the IM was not recovered.

Before moving the BF picomotors the position of the BF was

I used the BF picomotor 2 and some BF Damper coil magnet actuator to move it to

  • L, T, V = -738 , 51, 2303 um
  • R, P, Y = 472, -786, 5325 urad.

The position is not exaclty the same, especially in T and P, nevertheless, it's very similar. The IM alignment did not improve, the IM-Y magnitude is still large:

  • IM-L, T, V = 50 , 39, 60 um
  • IM-R, P, Y = -165, -676, -2658 urad.

See this notebook page for more details.

It's worth mentioning that the misalignment currently affecting the IM and IRM is large and the differences in  BF-T and P referred to above are relatively small and, likely, they don't play an important role in the problem.

The next step would be use the optic fibre camera to inspect the condition of the OSEM flags. One or more might have moved off from the magnets. By simple visual inpection they all look fine but they are in places where it's difficult to inspect and we might be having the wrong impression.

Search Help